OpenStudy (anonymous):

evaluate this integral by parts. (i think it's an easy one i just am drawing a blank)

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\tan^{-1}y dy \]

OpenStudy (mathmale):

Referring to the familiar Int(u*dv) = u*v - Int(v du) formula, why not let u=arctan y and du=dy? Have you tried this yet?

OpenStudy (mathmale):

If u = arctan y, then du = ( ? ) dy

OpenStudy (anonymous):

oooh. du=1/1+u^2 dy?

OpenStudy (mathmale):

Nice work. See what you can do from this point on; by all means message me if you'd like further discussion. Good luck!

OpenStudy (anonymous):

thank you! ah i havent remembered or looked at those derivatives in a long while. i think from here i am ok